# A Survey on Benchmarks of Multimodal Large Language Models

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) are gaining increasing popularity in both academia and industry due to their remarkable performance in various applications such as visual question answering, visual perception, understanding, and reasoning. Over the past few years, significant efforts have been made to examine MLLMs from multiple perspectives. This paper presents a comprehensive review of 180 benchmarks and evaluation for MLLMs, focusing on (1)perception and understanding, (2)cognition and reasoning, (3)specific domains, (4)key capabilities, and (5)other modalities. Finally, we discuss the limitations of the current evaluation methods for MLLMs and explore promising future directions. Our key argument is that evaluation should be regarded as a crucial discipline to better support the development of MLLMs. For more details, please visit our GitHub repository: https://github.com/swordlidev/Evaluation-Multimodal-LLMs-Survey.
@@ 1 Introduction
Large Language Models (LLM) have recently garnered substantial interest across academic and industrial domains. The impressive performance of LLMs such as GPT ouyang2022instructGPT has fueled optimism that they could represent a step towards Artificial General Intelligence (AGI) in this era. The remarkable abilities of LLM have inspired efforts to integrate them with other modality-based models to enhance multimodal competencies. Consequently, Multimodal Large Language Models (MLLMs) jin2024emllm have emerged, This concept is further supported by the extraordinary success of proprietary models like OpenAI’s GPT-4V achiam2023gpt4 and Google’s Geminiteam2023gemini . In contrast to earlier models that were limited to solving specific tasks, Multimodal Large Language Models (MLLMs) demonstrate exceptional performance across a variety of applications, including both general Visual Question Answering (VQA) tasks and domain-specific challenges.
A comprehensive and objective benchmark for evaluating MLLMs is essential for comparing and investigating the performance of various models, and it plays a crucial role in the success of MLLMs. First, evaluating LLMs helps us better understand the strengths and weaknesses of MLLMs. For example, the SEED-Bench li2023SEED-Bench illustrates that current MLLMs show weaker abilities in understanding spatial relationships between objects while achieving relatively high performance on global image comprehension. Second, evaluations across various scenarios can offer valuable guidance for MLLM applications in fields such as medicine, industry, and autonomous driving. This, in turn, can inspire future designs and expand the scope of their capabilities. Third, the broad applicability of MLLMs underscores the importance of ensuring their robustness, safety, and reliability, especially in safety-sensitive sectors. Finally, it is significant to evaluate other user-friendly features of MLLMs including the ability to handle long contexts and accurately follow instructions. Therefore, we aim to raise awareness in the community of the importance of MLLM evaluations by reviewing the current evaluation protocols.
@@
In this survey, we aim to provide a comprehensive overview of recent advancements in the rapidly evolving field of MLLM evaluation. As illustrated in Figure 1, our survey covers five key areas of MLLM evaluation, encompassing 20-30 detailed categories. The figure also illustrates the trend of MLLM evaluation papers over time, revealing a rapid increase in the number of publications. This growth indicates that the research area has garnered widespread attention. Furthermore, we provide statistics on the performance of the top three MLLMs across 83 benchmarks since 2024. The data indicates that OpenAI’s GPT-4 and Google’s Gemini exhibit superior performance and have attracted significant academic attention. as depicted in Figure. LABEL:fig:_MLLMs_evaluations_structure, we survey 180 benchmarks and organize the literature in a taxonomy consisting of five primary categories, encompassing various aspects, focusing on perception and understanding; cognition and reasoning; specific domains; key capabilities and other modalities.
@@ 2 Preliminaries
Figure. 1 compares several common MLLMs including GPT4achiam2023gpt4 , Geminiteam2023gemini , LLaVAliu2023llava , Qwen-VLbai2023Qwen-VL , Claudeanthropic2024claude , InstructBLIPdai2024instructblip , mPLUG-Owl2ye2023mplugowl2 , SPHINXlin2023SPHINX , Intern-VLchen2023internvl , Yi-VLai2024Yi-VL , VideoChat2li2023videochat , Video-LLaMAzhang2023Video-LLaMA , Cambrian-1tong2024Cambrian , PLLaVAxu2024PLLaVA , Blip2li2023blip2 , MiniGPT4-Videoataallah2024MiniGPT4-Video . The standard MLLM framework can be divided into three main modules: a visual encoder $g$ tasked with receiving and processing visual inputs, a pre-trained language model that manages the received multimodal signals and performs reasoning, and a visual-language projector $P$ which functions as a bridge to align the two modalities. A diagram of the architecture and training process is illustrated in Figure. 2. This figure outlines the base LLM, the vision encoder, the projector, as well as the pertaining and instruction tuning.
@@ 3.2 Fine-grained Perception
Visual Grounding and Object Detection: Object-level grounding and detection are critical steps in better perceiving images and solving image-related questions, providing a stronger link between QA pairs and images. Flickr30k Entities plummer2016Flickr30kEntities and Visual7W zhu2016Visual7W were early benchmarks for entity localization, focusing on the detailed grounding of specific phrases in image regions and QA tasks related to local areas. These benchmarks served as the foundation for subsequent developments in the field. However, they did not provide enough detailed contextual information about objects within the scene. Zang et al. zang2023CODE investigated contextual object detection, which involved understanding visible objects within human-AI interactive contexts, and presented the CODE benchmark to facilitate research in this area. To evaluate MLLMs’ ability in challenging scenarios where images contained abundant and complex information, Wu et al. wu2023V*Bench introduced V*Bench, a benchmark focused on detailed visual grounding in high-resolution images.
@@
Fine-grained Identification and Recognition:Fine-grained identification and recognition requires MLLMs to identify and analyze detailed visual features. In order to determine what makes for a good visual tokenizer, GVT-benchwang2023GVT-Bench was designed to evaluate MLLMs’ fine-grained visual perception through object counting and multi-class identification. However, the data used in GVT-bench were constrained to relatively small resolutions. To address this limitation, Li et al. li2023MagnifierBench extended the input resolution to a much larger range and introduced MagnifierBench, which tested MLLMs’ ability to discern the details of small objects in high-resolution images. Considering that most open-source MLLMs adopted the Pretrained Contrastive Language-Image PreTraining (CLIP) modelradford2021clip as the visual encoder, Tong et al. tong2024MMVP proposed the MMVP benchmark consisting of CLIP-blind pairs. They found that all tested models struggled with simple questions about visual details. Given that existing benchmarks were insufficient in size and did not cover crucial visual elements such as depth and spatial awareness, CV-Bench tong2024CV-Bench was created to evaluate the fundamental 2D and 3D visual understanding of MLLMs, using a significantly larger number of examples. Additionally, P2GB chen2024P2GB was designed to assess fine-grained visual capabilities, especially in text-rich scenarios, requiring comprehensive recognition and understanding of image text content. Moreover, VisualCoT shao2024VisualCoT was constructed with visual chain-of-thought prompts, allowing MLLMs to focus on specific regions within a complete image, similar to how humans comprehend intricate visual information.
@@ 4.1 General Reasoning
Visual Relation: Evaluating MLLMs’ reasoning of visual spatial relationships involves assessing their abilities to understand the spatial arrangement, relative positions, and interactions of objects. VSRVSR and What’s Upkamath2023what'sup were developed to test MLLMs’ ability to reason about spatial relations in natural image-text pairs. However, these benchmarks primarily focused on classification tasks rather than comprehending relations within scenes. To assess more complex and general relational reasoning of MLLMs, Wang et al.wang2024CRPE introduced the Circular-based Relation Probing Evaluation (CRPE), the first benchmark to encompass all elements of relation triplets (subject, predicate, object). Recognizing that interactions and associations between distinct objects remained a significant challenge for MLLMs, Nie et al.nie2024MMRel developed Multi-Modal Relation Understanding (MMRel), a large-scale benchmark consisting of data with well-defined inter-object relations. Additionally, GSR-BENCHrajabi2024GSR-BENCH extended the What’s Up benchmark by incorporating bounding box annotations and depth information to better evaluate grounded spatial relation understanding. For the evaluation of 3D spatial cognition in MLLMs, Chen et al.cheng2024SpatialRGPT proposed SpatialRGBT-Bench, a benchmark that incorporated ground-truth 3D annotations and flexibly integrated depth information.
@@ 5.1 Text-rich VQA
Chart-oriented Question Answering: Chart understanding plays a pivotal role when applying MLLMs to real-world tasks such as analyzing scientific papers or financial reports. ChartQAmasry2022ChartQA was a large-scale benchmark involving visual and logical reasoning over charts, but it had limited chart types. To scale up, Li et al.li2023SciGraphQA utilized Arxiv papers and constructed SciGraphQA with multi-turn question-answer conversations on scientific graphs. MMCliu2024MMC-Benchmark and ChartBenchxu2024ChartBench further expanded the types and tasks of chart data. However, the question design in these benchmarks remained relatively simplistic. To further validate MLLMs’ capabilities in more complex reasoning tasks involving chart data, Xia et al.xia2024ChartX proposed the benchmark ChartX for comprehensive chart understanding and reasoning. However, ChartX yielded artificial charts generated by GPT-4achiam2023gpt4 , leading to a narrow distribution. Therefore, CharXivwang2024CharXiv was proposed for evaluating MLLMs’ understanding of real-world scientific charts, including complex compositions with multiple subplots collected from arXiv. Moreover, in order to assess MLLMs’ understanding of charts from various perspectives, Fan et al.fan2024CHOPINLLM proposed CHOPINLLM, including three different levels of questions (literal, inferential, and reasoning QAs).
@@ 5.4 Other Applications
Medicine:Medical-based benchmarks primarily assess MLLMs’ ability to integrate medical knowledge with visual modalities to perform accurate medical diagnoses and recommendations. The Asclepius benchmarkwang2024Asclepius covered a range of medical specialties, aiming to comprehensively evaluate MLLMs’ capabilities across various medical fields. However, it primarily focused on 2D medical images, leaving 3D images less explored. To address this, Bai et al.bai2024M3D introduced M3D-Bench, a 3D multimodal medical benchmark that enabled automatic evaluation across eight tasks. Furthermore, Chen et al.chen2024GMAI-MMBench developed GMAI-MMBench, the most comprehensive general medical AI benchmark to date, featuring a well-categorized data structure and multi-perceptual granularity.
@@ 7.1 Videos
Comprehensive Evaluation: Comprehensive evaluation of video MLLMs focuses on their overall abilities, including video comprehension, scene recognition, temporal reasoning, and various other related tasks. With the development of video MLLMs, several works collected extensive video datasets to comprehensively evaluate MLLMs, such as Video-Bench ning2023Video-Bench and AutoEval-Video chen2024AutoEval-Video . However, they were limited by a lack of diversity in video types and insufficient coverage of temporal dynamics. To create a more comprehensive and high-quality assessment of MLLM in video, Fu et al. fu2024Video-MME proposed Video-MME, which offered more diverse video types, broader temporal coverage, and higher quality annotations. Additionally, due to the ability of videos to capture rich representations of real-world dynamics and causalities, some works evaluated MLLMs’ understanding of videos through the lens of "world models," which required MLLMs to interpret and reason about complex real-world dynamics. MMWorld he2024MMWorld and WorldNet ge2024WorldNet-Crafted were designed to rigorously evaluate the capabilities of MLLMs in world modeling via video understanding. MMWorld spanned a broad spectrum of disciplines, featuring a diverse array of question types for multi-faceted reasoning, while WorldNet encompassed millions of samples across a wide range of real-world scenarios and tasks.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) are gaining increasing popularity in both academia and industry due to their remarkable performance in various applications such as visual question answering, visual perception, understanding, and reasoning. Over the past few years, significant efforts have been made to examine MLLMs from multiple perspectives. This paper presents a comprehensive review of 200 benchmarks and evaluation for MLLMs, focusing on (1)perception and understanding, (2)cognition and reasoning, (3)specific domains, (4)key capabilities, and (5)other modalities. Finally, we discuss the limitations of the current evaluation methods for MLLMs and explore promising future directions. Our key argument is that evaluation should be regarded as a crucial discipline to support the development of MLLMs better. For more details, please visit our GitHub repository: https://github.com/swordlidev/Evaluation-Multimodal-LLMs-Survey.
@@ 1 Introduction
Large Language Models (LLM) have recently garnered substantial interest across academic and industrial domains. The impressive performance of LLMs such as GPT ouyang2022instructGPT has fueled optimism that they could represent a step towards Artificial General Intelligence (AGI) in this era. The remarkable abilities of LLM have inspired efforts to integrate them with other modality-based models to enhance multimodal competencies. Consequently, Multimodal Large Language Models (MLLMs) jin2024emllm have emerged, This concept is further supported by the extraordinary success of proprietary models like OpenAI’s GPT-4V achiam2023gpt4 and Google’s Geminiteam2023gemini . In contrast to earlier models that were limited to solving specific tasks, MLLMs demonstrate exceptional performance across a variety of applications, including both general Visual Question Answering (VQA) tasks and domain-specific challenges.
A comprehensive and objective benchmark for evaluating MLLMs is essential for comparing and investigating the performance of various models, and it plays a crucial role in the success of MLLMs. First, evaluating MLLMs helps us better understand the strengths and weaknesses of MLLMs. For example, the SEED-Bench li2023SEED-Bench illustrates that current MLLMs show weaker abilities in understanding spatial relationships between objects while achieving relatively high performance on global image comprehension. Second, evaluations across various scenarios can offer valuable guidance for MLLM applications in fields such as medicine, industry, and autonomous driving. This, in turn, can inspire future designs and expand the scope of their capabilities. Third, the broad applicability of MLLMs underscores the importance of ensuring their robustness, safety, and reliability, especially in safety-sensitive sectors. Finally, it is significant to evaluate other user-friendly features of MLLMs including the ability to handle long contexts and accurately follow instructions. Therefore, we aim to raise awareness in the community of the importance of MLLM evaluations by reviewing the current evaluation protocols.
@@
In this survey, we aim to provide a comprehensive overview of recent advancements in the rapidly evolving field of MLLM evaluation. As illustrated in Figure 1, our survey covers five key areas of MLLM evaluation, encompassing 20-30 detailed categories. The figure also illustrates the trend of MLLM evaluation papers over time, revealing a rapid increase in the number of publications. This growth indicates that the research area has garnered widespread attention. Furthermore, we provide statistics on the performance of the top three MLLMs across 83 benchmarks since 2024. The data indicates that OpenAI’s GPT-4 and Google’s Gemini exhibit superior performance and have attracted significant academic attention. As depicted in Figure. LABEL:fig:_MLLMs_evaluations_structure, we survey 200 benchmarks and organize the literature in a taxonomy consisting of five primary categories, encompassing various aspects, focusing on perception and understanding; cognition and reasoning; specific domains; key capabilities and other modalities.
@@ 2 Preliminaries
Figure. 1 compares several common MLLMs including GPT4achiam2023gpt4 , Geminiteam2023gemini , LLaVAliu2023llava , Qwen-VLbai2023Qwen-VL , Claudeanthropic2024claude , InstructBLIPdai2024instructblip , mPLUG-Owl2ye2023mplugowl2 , SPHINXlin2023SPHINX , Intern-VLchen2023internvl , Yi-VLai2024Yi-VL , VideoChat2li2023videochat , Video-LLaMAzhang2023Video-LLaMA , Cambrian-1tong2024Cambrian , PLLaVAxu2024PLLaVA , Blip2li2023blip2 , MiniGPT4-Videoataallah2024MiniGPT4-Video . The standard MLLM framework can be divided into three main modules: a visual encoder $g$ tasked with receiving and processing visual inputs, a pre-trained language model that manages the received multimodal signals and performs reasoning, and a visual-language projector $P$ which functions as a bridge to align the two modalities. A diagram of the architecture and training process is illustrated in Figure. 2. This figure outlines the base LLM, the vision encoder, and the projector, as well as the pretraining and instruction tuning.
@@ 3.2 Fine-grained Perception
Visual Grounding and Object Detection: Object-level grounding and detection are critical steps in better perceiving images and solving image-related questions, providing a stronger link between QA pairs and images. Flickr30k Entities plummer2016Flickr30kEntities and Visual7W zhu2016Visual7W were early benchmarks for entity localization, focusing on the detailed grounding of specific phrases in image regions and QA tasks related to local areas. These benchmarks served as the foundation for subsequent developments in the field. However, they did not provide enough detailed contextual information about objects within the scene. Zang et al. zang2023CODE investigated contextual object detection, which involved understanding visible objects within human-AI interactive contexts, and presented the CODE benchmark to facilitate research in this area. To evaluate MLLMs’ ability in challenging scenarios where images contained abundant and complex information, Wu et al. wu2023V*Bench introduced V*Bench, a benchmark focused on detailed visual grounding in high-resolution images. However, these benchmarks for visual grounding primarily used data with short captions and overlooked MLLMs’ chat performance when asked to ground. To address this issue, Zhang et al.zhang2023Grounding-Bench introduced a benchmark called Grounding-Bench, which allowed for combined evaluation of both grounding and chat capabilities.
@@
Fine-grained Identification and Recognition:Fine-grained identification and recognition requires MLLMs to identify and analyze detailed visual features. In order to determine what makes for a good visual tokenizer, GVT-benchwang2023GVT-Bench was designed to evaluate MLLMs’ fine-grained visual perception through object counting and multi-class identification. However, the data used in GVT-bench were constrained to relatively small resolutions. To address this limitation, MagnifierBenchli2023MagnifierBench and HR-Benchwang2024HR-Bench extended the input resolution to a significantly larger range, thereby evaluating the ability of MLLMs to discern details in high-resolution images. Considering that most open-source MLLMs adopted the Pretrained Contrastive Language-Image PreTraining (CLIP) modelradford2021clip as the visual encoder, Tong et al. tong2024MMVP proposed the MMVP benchmark consisting of CLIP-blind pairs. They found that all tested models struggled with simple questions about visual details. Given that existing benchmarks were insufficient in size and did not cover crucial visual elements such as depth and spatial awareness, CV-Benchtong2024CV-Bench was created to evaluate the fundamental 2D and 3D visual understanding of MLLMs. Furthermore, Yu et al.yu2024SPARK proposed a benchmark named SPARK, which focused on evaluating the ability of MLLMs to analyze images captured by various multi-vision sensors, including RGB, thermal, depth, and X-ray images. Additionally, P2GB chen2024P2GB was designed to assess fine-grained visual capabilities, especially in text-rich scenarios, requiring comprehensive recognition and understanding of image text content. Moreover, VisualCoT shao2024VisualCoT was constructed with visual chain-of-thought prompts, allowing MLLMs to focus on specific regions within a complete image, similar to how humans comprehend intricate visual information.
@@ 4.1 General Reasoning
Visual Relation: Evaluating MLLMs’ reasoning about visual relations primarily involves spatial and causal relationships. Regarding spatial relationships, these benchmarks mainly focus on evaluating MLLMs’ understanding of the spatial arrangement, relative positions, and interactions of objects jin2024llava-vsd . VSRVSR and What’s Upkamath2023what'sup were developed to test MLLMs’ ability to reason about spatial relations in natural image-text pairs. However, these benchmarks primarily focused on classification tasks rather than comprehending relations within scenes. To assess more complex and general relational reasoning of MLLMs, Wang et al.wang2024CRPE introduced the Circular-based Relation Probing Evaluation (CRPE), the first benchmark to encompass all elements of relation triplets (subject, predicate, object). Recognizing that interactions and associations between distinct objects remained a significant challenge for MLLMs, Nie et al.nie2024MMRel developed Multi-Modal Relation Understanding (MMRel), a large-scale benchmark consisting of data with well-defined inter-object relations. Additionally, GSR-BENCHrajabi2024GSR-BENCH extended the What’s Up benchmark by incorporating bounding box annotations and depth information to better evaluate grounded spatial relation understanding. Similarly, Chen et al.cheng2024SpatialRGPT proposed SpatialRGBT-Bench, a benchmark that incorporated ground-truth 3D annotations and flexibly integrated depth information. In terms of causal relationships, causal reasoning refers to identifying the relationship between a cause and its effect. To explore MLLMs’ ability to perform multimodal causal reasoning tasks, Li et al.li2024MuCR proposed a benchmark called MuCR. This benchmark challenged MLLMs to infer semantic cause-and-effect relationships based solely on visual cues such as action, appearance, clothing, and environment.
@@
Chain-of-Thought Reasoning:Multi-modal Chain-of-Thought (CoT) requires models to leverage knowledge from both textual and visual modalities for step-by-step reasoning. To diagnose the multi-hop reasoning ability and interpretability of multi-modal large language models (MLLMs), Lu et al.lu2022ScienceQA generated lectures and explanations as CoT to mimic the multi-hop reasoning process for answering questions and presented Science Question Answering (SCIENCEQA). SCIENCEQA consisted of questions covering diverse science topics, along with annotations of their answers and corresponding lectures and explanations, demonstrating the utility of CoT in improving the question-answering performance of MLLMs. Besides, Shao et al.shao2024VisualCoT introduced the visual chain-of-thought benchmark for evaluating MLLMs in scenarios where they need to focus on specific local regions or reasons to identify objects. However, these benchmarks were relatively simple and focused only on certain specific domains. Therefore, Chen et al.chen2024M3CoT introduced M3CoT to address these limits, advancing the multi-domain, multi-step, and multi-modal CoT.
@@ 5.1 Text-rich VQA
Chart-oriented Question Answering: Chart understanding plays a pivotal role when applying MLLMs to real-world tasks such as analyzing scientific papers or financial reports. ChartQAmasry2022ChartQA was a large-scale benchmark involving visual and logical reasoning over charts, but it had limited chart types. To scale up, Li et al.li2023SciGraphQA utilized Arxiv papers and constructed SciGraphQA with multi-turn question-answer conversations on scientific graphs. MMCliu2024MMC-Benchmark and ChartBenchxu2024ChartBench further expanded the types and tasks of chart data. However, the question design in these benchmarks remained relatively simplistic. To further validate MLLMs’ capabilities in more complex reasoning tasks involving chart data, Xia et al.xia2024ChartX proposed the benchmark ChartX for comprehensive chart understanding and reasoning. However, ChartX yielded artificial charts generated by GPT-4achiam2023gpt4 , leading to a narrow distribution. Therefore, CharXivwang2024CharXiv was proposed for evaluating MLLMs’ understanding of real-world scientific charts, including complex compositions with multiple subplots collected from arXiv. To further improve the quality, Roberts et al.roberts2024SciFIBench presented SciFIBench, a benchmark for scientific figure interpretation. They adopted adversarial filtering when curating negative examples for each question to increase difficulty, while also performing human verification on every question to ensure high-quality content. Moreover, in order to assess MLLMs’ understanding of charts from various perspectives, Fan et al.fan2024CHOPINLLM proposed CHOPINLLM, including three different levels of questions (literal, inferential, and reasoning QAs).
@@ 5.4 Other Applications
Medicine:Medical-based benchmarks primarily assess MLLMs’ ability to integrate medical knowledge with visual modalities to perform accurate medical diagnoses and recommendations. The Asclepius benchmarkwang2024Asclepius covered a range of medical specialties, aiming to comprehensively evaluate MLLMs’ capabilities across various medical fields. However, it primarily focused on 2D medical images, leaving 3D images less explored. Therefore, Bai et al.bai2024M3D introduced M3D-Bench, a 3D multimodal medical benchmark that enabled automatic evaluation across eight tasks. Furthermore, Chen et al.chen2024GMAI-MMBench developed GMAI-MMBench, the most comprehensive general medical AI benchmark to date, featuring a well-categorized data structure and multi-perceptual granularity. However, these benchmarks were relatively lacking in depth, such as overlooking the rich interconnected information across many heterogeneous biomedical sensors. To address this challenge, To address this challenge, Mo et al.mo2024MultiMed presented MultiMed, a benchmark designed to evaluate the ability of MLLMs to integrate data from multiple sources while performing several tasks simultaneously.
@@ 7.1 Videos
Comprehensive Evaluation: Comprehensive evaluation of video MLLMs focuses on their overall abilities, including video comprehension, scene recognition, temporal reasoning, and various other related tasks. With the development of video MLLMs, several works collected extensive video datasets to comprehensively evaluate MLLMs, such as Video-Bench ning2023Video-Bench and AutoEval-Video chen2024AutoEval-Video . However, they were limited by a lack of diversity in video types and insufficient coverage of temporal dynamics. To create a more comprehensive and high-quality assessment of MLLM in video, Fu et al.fu2024Video-MME proposed Video-MME, which offered more diverse video types, broader temporal coverage, and higher quality annotations. Nevertheless, Video-MME employed a multiple-choice question-and-answer format for straightforward and flexible assessment, but this format overlooked the quality and richness of free-form expressions generated by MLLMs. To address this issue, Fang et al.fang2024MMBench-Video introduced MMBench-Video, a comprehensive benchmark that consisted of free-form questions mirroring practical use cases. Additionally, due to the ability of videos to capture rich representations of real-world dynamics and causalities, some works evaluated MLLMs’ understanding of videos through the lens of "world models," which required MLLMs to interpret and reason about complex real-world dynamics. MMWorld he2024MMWorld and WorldNet ge2024WorldNet-Crafted were designed to rigorously evaluate the capabilities of MLLMs in world modeling via video understanding. MMWorld spanned a broad spectrum of disciplines, featuring a diverse array of question types for multi-faceted reasoning, while WorldNet encompassed millions of samples across a wide range of real-world scenarios and tasks.